A paint manufacturer uses a machine to fill gallon cans with paint. The manufacturer wants to estimate the mean volume of paint of the machine is putting in the cans within 0.3 ounce. Determine the minimum sample size required to construct a 89% confidence interval for the population mean. Assume the population standard deviation is 0.8 ounce.
1
Expert's answer
2020-12-28T11:20:59-0500

"n=(\frac{z\sigma}{0.3})^2"

"Z_{0.055}=1.5982"

"n=(\frac{1.5982×0.8}{0.3})^2=" 18.16

n=18.